OpenStudy (anonymous):

What is the value of f(x) = 36^x when x = 1/2?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

f(1/2)=36^(1/2) means the square root of 36 f(1/2)= 6

OpenStudy (anonymous):

Thank you.